Mr. Amos invests $1,500 today with the hope that in 7 years her investment will be worth $23,500. The investment will compound semiannually. At the end of 7 years, what will be Hanna’s annual rate of return?
Answer #1

PV =1500
Number of Periods =7*2=14
FV =23500
FV=PV*(1+r/2)^14
23500=1500*(1+r/2)^14
(1+r/2)^14 =23500/1500
(1+r/2)^14 =15.67
r /2 =(15.67^(1/14)-1)
Hanna's annual rate of return =43.44%